# Cascade switching current detectors based on arrays of Josephson junctions

### v2

@@ Results
Fig. 2 (d) shows the $I$-$V$s of another mesa at different MW powers and $I_{b}=38.1~{}\mu$A. The cascade gain depends on bias. Fig. 2 (e) represents the MW power dependencies of $R$ (symbols) and $n$ (lines), measured at two biases. At lower $I_{b}=25.3\ \upmu\mathrm{A}$, the sensitivity, $dR/dP$, remains small up to some threshold, $P\sim 0.2~{}P_{max}$. At higher $I_{b}=38.1~{}\upmu\mathrm{A}$, there is no threshold and the sensitivity is high at $P=0$. With further increase of $I_{b}$, the sensitivity remains high at $P=0$ but saturates at successively smaller $P$, thus reducing the dynamic range [8]. Therefore, there is an optimal bias ($I_{b}=38.1\ \upmu\mathrm{A}$ in this case) with both high sensitivity and large dynamic range. For comparison, in the Supplementary we analyze the ordinary SCD operation on the same mesa using a single surface IJJ. The surface IJJ has much smaller $I_{s}$ [27] and can be measured without activation of other IJJs [19]. The maximum single-junction response is $\Delta R\sim 40~{}\Omega$, much smaller than $\sim 8~{}\mathrm{k}\Omega$ for the cascade SCD in Fig. 2 (e) at the same MW power for the same mesa.
@@
Fig. 2 (f) shows the MW directivity of this mesa. It exhibits a four-fold shape. The two smaller lobes are aligned with the whisker (indicated by the blue line). The larger lobes are approximately perpendicular, slightly inclined in the direction of one of bias electrodes (see the Supplementary for more details). The large-scale electrode geometry of this device is shown in the background (yellow). The observed four-fold directivity is consistent with the whisker-mesa-electrode geometry, which form a turnstile-antenna, as shown in Fig. 2 (b) [50].
@@
To verify THz operation, we perform an in-situ generation-detection experiment in which one Bi-2212 mesa emits Josephson radiation and a nearby mesa detects it [19, 22]. Fig. 3 (a) shows the $I$-$V$ of a generator mesa ($\sim 10\times 15~{}\mu$m2), containing $N=160\pm 10$ IJJs. We analyze the downturn part of it (magenta) at which all $N$ IJJs are active. In this case, the Josephson frequency is well defined, $f_{J}=2eV/hN$, as indicated in the top axis. Emission from Bi-2212 mesas occurs at cavity modes inside the mesa [19, 32, 33]. The key signature of such emission (as opposed to heating) is the non-monotonous dependence on the dissipation power, $P_{\text{gen}}=I_{\text{gen}}V_{\text{gen}}$, with distinct peaks at cavity resonances [19].
Fig. 3 (b) shows a low-bias part of the $I$-$V$ of a detector mesa ($\sim 15\times 15~{}\mu$m2), about $22~{}\mu$m away from the generator mesa. Fig. 3 (c) is a close-up on the central part, representing the $I$-$V$ of the surface junction. It has $I_{c}^{\prime}\simeq 15~{}\mu$A, significantly smaller than $I_{c}\simeq 370~{}\mu$A for the rest of the IJJs. This disparity allows ordinary SCD measurements at $I_{c}^{\prime}<I_{b}\ll I_{c}$, which activates only one surface IJJ [19]. The cascade SCD can be analyzed on the same mesa by increasing bias to $I_{b}\gtrsim I_{c}$.
Fig. 3 (d) shows single-junction (olive) and cascade (magenta) SCD responses versus $P_{gen}$, measured at bias amplitudes, $I_{b}=25$ and 370 $\mu$A, respectively. It can be seen that (in this case) the resolution of a single-junction SCD is not sufficient for resolving the emission. The cascade SCD demonstrates a much larger response with clearly visible emission peaks. The main emission for this generator mesa occurs in the range from $\sim 4$ to 8.4 THz, as indicated by the top axis in Fig. 3 (d). Fig. 3 (e) shows the detector $I$-$V$’s measured with $I_{b}=370~{}\mu$A at several bias points of the generator, marked by the same-color circles in Figs. 3 (a) and (d). It can be seen that cascade amplification factor, $n$, is playing the dominating role in the detector response.
@@ Discussion
We observed similar behavior both for low- and high-$T_{c}$ devices, either linear arrays or stacks. Both types of detectors have some advantages. 2D-linear arrays facilitate a large absorption area and a simple scalability to tens of thousand of JJs [44].
The list of advantages of high-$T_{c}$ devices include: (i) Broad frequency range covering the whole THz area [0.1-10] THz [19]. (ii) Compactness in combination with large $N$. (iii) Large $V_{c}\sim 20-30$ mV [26, 27], enabling high readout signals. From Figs. 1 (b) and 2 (c) it can be seen that the maximum readout voltage, $NV_{c}$, for a high-$T_{c}$ mesa (${}5$V) is more than 100 times larger than for a low-$T_{c}$ array ($\sim 0.04$ V) despite only a factor two difference in $N$ (250 vs. 128). (iv) The atomic separation of IJJs enables the strongest interjunction coupling, needed for the avalanche switching.
@@
[…]
@@
The detector sensitivity $S$(V/W) can be calculated in the following way. The readout voltage from the array is $V(n)=nV_{c}$. The spread of switching currents for JJs in the array can be written as, $I_{s}(n)\simeq I_{s}^{*}+\sigma(n/N-1/2)$, where $I_{s}^{*}$ is the mean switching current and $\sigma$ is the width of the switching current distribution. At the optimal bias, $I_{b}\simeq I_{s}^{*}$, the effective differential resistance is, $R_{d}\simeq NV_{c}/\sigma$. The response of SCD is $\Delta V\simeq R_{d}I_{THz}$, where $I_{THz}$ is the induced THz current. The absorbed THz power can be writen as, $P_{THz}=I_{THz}^{2}R_{THz}/2$, where $R_{THz}$ is the real part of the high-frequency impedance of the detector (including electrodes). Thus, we obtain
$S\simeq\frac{2NV_{c}}{\sigma R_{THz}I_{THz}}.$ (1)
@@
The procedure for quantitative estimation of sensitivity and noise-equivalent power (NEP) is described in Methods. For the case of Bi-2212 detectors from Figs. 2 and 3 we get $S>10^{7}$ (V/W) and $NEP\sim 10^{-17}~{}$(W$/\sqrt{Hz}$) for both MW and THz detection. There are very good numbers, however, far from optimal. Lets estimate achievable characteristics for a small, high-quality Bi-2212 mesa, like in Fig. 2 (c). Assuming $N=500$, $V_{c}=20$, $I_{s}^{*}=50~{}\upmu$A; $\sigma=0.01I_{s}^{*}$ (typical for fluctuation-induced $\sigma$ at $T\sim 4$ K [49, 8]); $R_{THz}=100~{}\Omega$ (for an efficient detector $R_{THz}$ should be comparable to the impedance of free space); and $I_{THz}=0.1\sigma=50~{}$nA; we obtain $S\simeq 8\cdot 10^{12}$ (V/W) and $NEP\simeq 10^{-22}~{}$(W$/\sqrt{Hz}$). This estimation is made under realistic assumptions and does not represent the ultimate limit of SCD [8].
@@
Two factors contribute to this high sensitivity: the large cascade gain $\propto N$ and the narrow switching width $\sigma$. The latter depends on the uniformity of the array, inter-junction coupling, thermal fluctuations and electrical noise. When all junction in the stack are identical, they will switch at the same bias current. A spread of parameters cause broadening of $\sigma$. However, coupling between JJs leads to current-locking even of slightly dissimilar junctions [36]. The true phase-locking of JJs can further reduce $\sigma$ [51]. Therefore, a strong inter-junction coupling is crucial for such a detector. Ultimately, $\sigma$ is limited by quantum fluctuations [14, 16, 17, 18, 24, 49], qualifying SCD as a quantum-limited photon detector.
A good detector should possess both high sensitivity to absorbed power and high absorption efficiency of incident radiation. The maximum efficiency ($50\%$) is achieved at the impedance matching condition [41, 40, 39]. However, a single JJ alone cannot provide proper matching because it is much smaller than $\lambda_{0}$ [39]. Therefore, the geometry of electrodes forming a receiving antenna becomes crucial. In the case of linear Nb arrays, long electrodes function as traveling wave antennas, as evidenced by off-axis directivities of both emission [45, 46] and absorption, Fig. 1 (f). Whisker-based Bi-2212 mesas have a turnstile antenna geometry [35], Fig. 2(b), enabling high radiation efficiency ($12\%$ at $f\simeq 4$ THz) [22] and causing the four-fold receiving directivity, Fig. 2 (f) [50]. Thus, the analysis of directivity (geometry-specific antenna effects) provides a valuable tool for investigating impedance matching and the efficiency of both absorption and emission.
@@
In conclusion, we have described the operational principle of a cascade switching current detector, utilizing arrays of coupled Josephson junctions. Coupling between junctions leads to the current locking phenomenon and results in avalanche-like switching of many junctions. This amplifies the readout voltage and enhances detector sensitivity. We have demonstrated two prototypes based either on low-$T_{c}$ linear Nb arrays, or high-$T_{c}$ stacked intrinsic Josephson junctions. Good impedance matching and absorption efficiency of the studied devices is confirmed by observation of geometry-specific antenna effects in receiving directivities. We conclude that Bi-2212 whisker-based cascade SCD are promising candidates for broad-band and high-performance THz sensors.
@@ Methods
The absorbed power, $P_{a}$, can be estimated from the suppression of the switching current, $\Delta I_{s}$, caused by the generated high-frequency current. In the MW case (low frequency limit) $\Delta I_{s}$ is just equal to the induced MW current [8]. $P_{a}$ is calculated using Eq. (4) from Ref. [19]: $P_{a}\simeq(2\sqrt{2}/3\pi)(\Delta I_{s}/I_{c0})^{3/2}I_{c0}V_{c}$. For the case of Fig. 2 (d), $\Delta I_{s}\simeq 3~{}\mu$A at the attenuation 3.7 dB, $I_{c0}=38~{}\mu$A, $V_{c}=20~{}$mV, yielding $P_{a}\simeq 5$ nW, approximately 30% of $P_{MW}$. The maximum voltage is $V\simeq 0.37$ V, and the sensitivity with respect to the absorbed power, $S_{a}=V/P_{a}\simeq 7.4~{}10^{7}$ (V/W) and the net sensitivity $S_{MW}=V/P_{MW}\simeq 2.6~{}10^{7}$ (V/W).
@@
The noise-equivalent power is determined as $NEP=v_{n}/S$, where $v_{n}$ is the voltage noise. It is typically in the range of few nV$/\sqrt{Hz}$ at room temperature, but for lock-in measurements at low $T$ it is certainly below 1 nV. Taking $v_{n}=1~{}$nV$/\sqrt{Hz}$ we obtain the following values of NEP with respect to absorbed power: $NEP_{MW}\simeq 1.35~{}10^{-17}$ W$/\sqrt{Hz}$ and $NEP_{THz}\simeq 6.25~{}10^{-17}$ W$/\sqrt{Hz}$.
The achieved characteristics are very good, but far from optimal, as discussed above. The main reason for the larger $NEP_{THz}$ than $NEP_{MW}$ is that the mesa used for the THz test has a factor 10 larger $I_{c0}$. The NEP of SCD is approximately proportional to the Josephson energy, $E_{J}=(\Phi_{0}/2\pi)I_{c0}$ [8]. Therefore, the weighted (by $E_{J}$) THz sensitivity is actually twice better than for MW.

### v3

@@ Introduction
Fig. 5 (d) shows the $I$-$V$s of the third mesa at different MW powers and $I_{b}=38.1~{}\mu$A. The cascade gain depends on $I_{b}$. Fig. 5 (e) represents the MW power dependencies of $V$ (symbols) and $n$ (lines), measured at two biases. At lower $I_{b}=25.3\ \upmu\mathrm{A}$, the sensitivity, $dV/dP$, remains small up to some threshold, $P\sim 0.2~{}P_{max}$. At higher $I_{b}=38.1~{}\upmu\mathrm{A}$, there is no threshold and the sensitivity is high at $P=0$. With further increase of $I_{b}$, the sensitivity remains high at $P=0$ but saturates at successively smaller $P$, thus reducing the dynamic range [15]. Therefore, there is an optimal bias ($I_{b}=38.1\ \upmu\mathrm{A}$ in this case) with both high sensitivity and large dynamic range. For comparison, in the Supplementary we analyze the ordinary SCD operation on the same mesa using a single surface IJJ. The surface IJJ has much smaller $I_{s}$ [37] and can be measured without activation of other IJJs [29]. The maximum single-junction response is almost 1000 times smaller than for the cascade SCD in Fig. 5 (e) at the same MW power for the same mesa.
Figure 6: Detection of THz radiation from a Bi-2212 mesa. (a) The $I$-$V$ of the generator mesa for upward (blue) and downward (magenta) bias sweeps. Top axis indicates the anticipated Josephson frequency. (b) and (c) Low-bias $I$-$V$’s of the detector mesa with (c) showing the $I$-$V$ of the weak surface junction. (d) Measured detector responses versus $P_{gen}=I_{gen}V_{gen}$. The olive line represents a single surface IJJ response (multiplied by a factor 50). The magenta curve represents the cascade SCD response of the same mesa. (e) A set of detector $I$-$V$’s with $I_{b}=370~{}\mu$A at different bias points of the generator, marked by the same-color circles in (a) and (d).
@@
Fig. 6 (f) shows the polarization-loss diagram of this mesa. It has a four-fold shape with two smaller lobes aligned with the whisker (indicated by the blue line) and the larger lobes approximately perpendicular, slightly inclined in the direction of one of bias electrodes (see the Supplementary for more details). The large-scale electrode geometry of this device is shown in the background (yellow). The observed four-fold diagram is consistent with the turnstile-antenna geometry of the device, as sketched in the inset [59].
@@
To verify THz operation, we perform an in-situ generation-detection experiment in which one mesa emits Josephson radiation and a nearby mesa detects it [29, 32]. Fig. 6 (a) shows the $I$-$V$ of a generator mesa ($\sim 10\times 15~{}\mu$m2), containing $N=160\pm 10$ IJJs. We analyze the downturn part of it (magenta) at which all $N$ IJJs are active. In this case, the Josephson frequency is well defined, $f_{J}=2eV/hN$, as indicated in the top axis. The emission occurs at cavity modes in the mesa [29, 42, 43]. The key signature of such emission (as opposed to heating) is the non-monotonous dependence of the signal on the dissipation power, $P_{\text{gen}}=I_{\text{gen}}V_{\text{gen}}$, with distinct peaks at cavity resonances [29].
Fig. 6 (b) shows a low-bias part of the $I$-$V$ of a detector mesa ($\sim 15\times 15~{}\mu$m2), about $22~{}\mu$m away from the generator mesa. Fig. 6 (c) is a close-up on the central part, representing the $I$-$V$ of the surface junction. It has $I_{c}^{\prime}\simeq 15~{}\mu$A, significantly smaller than $I_{c}\simeq 370~{}\mu$A for the rest of the IJJs. This disparity allows ordinary SCD measurements at $I_{c}^{\prime}<I_{b}\ll I_{c}$, which activates only one surface IJJ [29]. The cascade SCD can be analyzed on the same mesa by increasing bias to $I_{b}\sim I_{c}$.
Fig. 6 (d) shows single-junction (olive, ×$\times 50$) and cascade (magenta) SCD responses versus $P_{gen}$, measured at bias amplitudes, $I_{b}=25$ and 370 $\mu$A, respectively. It can be seen that (in this case) the resolution of a single-junction SCD is not sufficient for resolving the emission. The cascade SCD demonstrates a much larger response with clearly visible emission peaks. The main emission for this generator mesa occurs in the range from $\sim 4$ to 8.4 THz, as indicated in the top axis. Fig. 6 (e) shows the detector $I$-$V$’s measured with $I_{b}=370~{}\mu$A at several bias points of the generator, marked by the same-color circles in Figs. 6 (a) and (d). It can be seen that cascade amplification factor, $n$, is playing the dominating role in the detector response.
@@ Discussion
We observed similar behavior both for low- and high-$T_{c}$ devices, either linear arrays or stacks. Both types of detectors have some advantages. 2D-linear arrays facilitate a large absorption area and simple scalability to tens of thousands of JJs [52].
The list of advantages of high-$T_{c}$ devices include: (i) Broad frequency range covering the whole THz area [0.1-10] THz [29]. (ii) Compactness in combination with large $N$. (iii) Large $V_{c}\sim 20-30$ mV [36, 37], enabling high readout signals. From Figs. 4 (a) and 5 (b) it can be seen that the maximum readout voltage, $NV_{c}$, for a high-$T_{c}$ mesa (${}5$V) is more than 100 times larger than for the low-$T_{c}$ array ($\sim 0.04$ V) despite only a factor two difference in $N$ (250 vs. 128). (iv) The atomic separation of IJJs enables the strongest interjunction coupling, needed for the avalanche switching.
@@
As described in the Supplementary, for Bi-2212 detector, the signal, $\Delta V$, increases from 0.32 mV for the single junction SCD (no amplification, $n=1$), to 81 mV at sub-optimal bias, and almost up to 0.3 V at the optimal bias, Fig. 5 (e). We emphasize that these measurements were performed under the same conditions with the only difference in $I_{b}$, determining the cascade amplification.
@@
The sensitivity $S$(V/W) can be calculated in the following way. The readout voltage from the array is $V(n)=nV_{c}$. The spread of switching currents for JJs in the array can be written as, $I_{s}(n)\simeq I_{s}^{*}+\sigma_{n}(n/N-1/2)$, where $I_{s}^{*}$ is the mean switching current and $\sigma_{n}$ is the width of the multi-junction switching current distribution. At the optimal bias, $I_{b}\simeq I_{s}^{*}$, the SCD response is $\Delta V\sim NV_{c}(I_{THz}/\sigma_{n})$. Using Eq.(3), we obtain
$S_{n}\simeq\frac{2\chi NV_{c}}{\sigma_{n}R_{THz}I_{THz}}.$ (7)
@@
The procedure for quantitative estimation of sensitivity and noise-equivalent power (NEP) is described in Methods. For the case of Bi-2212 detectors from Figs. 5 and 6 we get $S>10^{7}$ (V/W) and $NEP\sim 10^{-17}~{}$(W$/\sqrt{Hz}$) for both MW and THz detection. There are very good numbers, however, far from optimal. Lets estimate achievable characteristics for a high-quality Bi-2212 mesa, like in Fig. 5 (b). Assuming $N=500$, $V_{c}=20$ mV, $I_{s}^{*}=50~{}\upmu$A; $\sigma_{n}=0.01I_{s}^{*}$ (see Fig. 5 (c) at $T\sim 4$ K) [58, 15]); $R_{THz}=100~{}\Omega$ (for an efficient detector $R_{THz}$ should be comparable to the impedance of free space); and $I_{THz}=0.1\sigma_{n}=50~{}$nA; we obtain $S\simeq 8\cdot 10^{12}$ (V/W) and $NEP\simeq 0.1$ (zW$/\sqrt{Hz}$). This estimation is made under realistic assumptions and does not represent the ultimate limit of SCD [15].
@@
Two factors contribute to this high sensitivity: the large cascade gain $\propto N$ and the narrow switching width $\sigma_{n}$. The latter depends on the uniformity of the array, inter-junction coupling, thermal fluctuations and electrical noise. When all junction in the stack are identical, they will switch at the same bias current. A spread of parameters cause broadening of $\sigma_{n}$. However, coupling between JJs leads to current-locking even of slightly dissimilar junctions [45]. The true phase-locking of JJs can further reduce $\sigma_{n}$ [60]. Therefore, a strong inter-junction coupling is crucial for such a detector. Ultimately, $\sigma_{n}$ is limited by quantum fluctuations [26, 22, 28, 4, 23, 34, 58], qualifying SCD as a quantum-limited photon detector.
@@
In conclusion, we have described the operational principle of a cascade switching current detector, utilizing arrays of coupled Josephson junctions. The coupling between junctions leads to the current locking phenomenon and results in avalanche-like switching of many junctions. This amplifies the readout voltage without changing the noise background and, therefore, enhances the signal-to-noise ratio and reduces NEP. We have demonstrated two prototypes based either on low-$T_{c}$ linear Nb arrays, or high-$T_{c}$ stacked intrinsic Josephson junctions. Good impedance matching and absorption efficiency of the studied devices is confirmed by observation of geometry-specific antenna effects. We conclude that Bi-2212 whisker-based cascade SCD are promising candidates for broad-band and high-performance THz sensors.
@@ Methods
Experimental details
Measurements are performed in a closed-cycle optical cryostat with a base temperature $\sim 3$ K at zero magnetic field. Arrays are current-biased using a programmable voltage source. The $I$-$V$ characteristics are measured in a quasi 4-probe configuration. Additional information about experimental setup, bias configuration and microwave measurements can be found in sec. SVII of the Supplementary. Details about fabrication and characterization of Nb arrays and Bi-2212 mesas are provided in sec. SVIII and SIX.
The operation of SCD requires statistical analysis. The switching statistics of Bi-2212 mesas for a single and multi-junction switching can be found in sec. SIV of the Supplementary. It is well described by conventional thermal-activation theory, described in sec. SII.
The SCD response is measured via lock-in measurements. For this, a harmonic current, $I=I_{b}\sin(2\pi f_{b}t)$, with the amplitude $I_{b}$ and frequency $f_{b}=23$ Hz is supplied for a time interval of 1 s. The read-out voltage corresponds to the Fourier component of the detector $V(t)$ wave form at $f=f_{b}$. The lock-in response is calculated on-flight using FPGA-based FFT routine. Additional information about SCD operation and lock-in readout can be found in sec. SI and SV of the Supplementary.
@@
The absorbed power, $P_{a}$, can be estimated from the suppression of the switching current, $\Delta I_{s}$, caused by the generated high-frequency current. In the MW case (low frequency limit) $\Delta I_{s}$ is just equal to $I_{MW}$ [15]. $P_{a}$ is calculated as [29]: $P_{a}\simeq(2\sqrt{2}/3\pi)(\Delta I_{s}/I_{c0})^{3/2}I_{c0}V_{c}$. For the case of Fig. 5 (d), $\Delta I_{s}\simeq 3~{}\mu$A at the attenuation 3.7 dB, $I_{c0}=38~{}\mu$A, $V_{c}=20~{}$mV, yielding $P_{a}\simeq 5$ nW, approximately 30% of $P_{MW}$. The maximum voltage is $V\simeq 0.37$ V, and the sensitivity with respect to the absorbed power, $S_{a}=V/P_{a}\simeq 7.4~{}10^{7}$ (V/W) and the net sensitivity $S_{MW}=V/P_{MW}\simeq 2.6~{}10^{7}$ (V/W).
@@
The noise-equivalent power is determined as $NEP=v_{n}/S$, where $v_{n}$ is the voltage noise. As shown in the Supplementary, in our setup it is in the range of few nV$/\sqrt{Hz}$, but could be reduced to below 1 nV$/\sqrt{Hz}$ using cryogenic preamplifiers. Taking $v_{n}=1~{}$nV$/\sqrt{Hz}$ we obtain the following values of NEP with respect to absorbed power: $NEP_{MW}\simeq 1.35~{}10^{-17}$ W$/\sqrt{Hz}$ and $NEP_{THz}\simeq 6.25~{}10^{-17}$ W$/\sqrt{Hz}$.
The achieved characteristics are very good, but far from optimal, as discussed above. The main reason for the larger $NEP_{THz}$ than $NEP_{MW}$ is that the mesa used for the THz test has a factor 10 larger $I_{c0}$. The NEP of SCD is approximately proportional to the Josephson energy, $E_{J0}$ [15]. Therefore, the weighted (by $E_{J0}$) THz sensitivity is actually twice better than for MW.
Signal-to-noise ratio
The signal-to-noise ratio in a cascade SCD is determined by the three quantities:
- The mean readout voltage $V_{n}$,
- The standard deviation (statistical uncertainty) $\delta V_{n}$,
- The noise threshold of readout electronics.
Direct noise measurements in our setup are presented in sec. SVI of the Supplementary. The noise floor is in the range 1-10 nV/Hz1/2, limited by the intrinsic noise in the room temperature preamplifier.
The effect of statistical uncertainty and the difference in sensitivities of single junction SCD and cascade SCD are discussed in sec. SIII. There we also write explicit equations for the sensitivity and provide corresponding numerical simulations.
The bottom line of this analysis is the following: Cascading increases the mean readout $V_{n}$, generally improves the relative uncertainty, $\delta V_{n}/V_{n}$, and usually expands the dynamic range. But it does not affect the measurement noise (determined by the preamplifier). Therefore, cascading always increases the signal-to-noise ratio and reduces NEP, compared to a single-junction SCD, just as sketched in Fig. 1 (a).
A good detector should have a high absorption efficiency of incident radiation. The maximum efficiency ($50\%$) is achieved at the impedance matching condition [24, 25]. However, a single JJ alone cannot provide proper matching because it is much smaller than $\lambda_{0}$ [25]. Therefore, the geometry of electrodes forming a receiving antenna, becomes crucial. In the case of linear Nb arrays, long electrodes function as traveling wave antennas, as evidenced by off-axis diagrams of both emission [53, 54] and absorption, Fig. 4 (c). Bi-2212 whisker mesas have a turnstile antenna geometry, enabling good impedance matching. This is evidenced by the high radiation efficiency ($12\%$ at $f\simeq 4$ THz) [32] and the four-fold polarization diagram reported in Fig. 5 (f) [59]. Thus, analysis of geometry-specific antenna effects provides a valuable tool for investigating impedance matching and the efficiency of both absorption and emission.